OpenStudy (melissa_something):

Easy fan and metal :) Write the standard form of the equation of the circle with the given center and radius:

OpenStudy (melissa_something):

(7,4); \[\sqrt{19}\]

OpenStudy (bibby):

if it's easy, why can't you do it?

OpenStudy (melissa_something):

Because it has a freaking radical!! and I wanna check my work @bibby

OpenStudy (bibby):

radicals are just things raised the 1/2 power post your work first then

OpenStudy (bibby):

recall that a circle of the form \((x-h)^2+(y-j)^2=r^2\) has center (h,j) with radius r also recall that squaring a square root cancels it. \((\sqrt3)^2=3\)

OpenStudy (melissa_something):

isn't it h and k?..

OpenStudy (bibby):

they're arbitrary letters. it's like when I call my wife tanya (her name is Rachel). She gets mad about it but names are names

OpenStudy (melissa_something):

Well I only plugged it in (x-7)^2 +(y-4)^2 = sqrt19

OpenStudy (melissa_something):

Lmao tanya's way off

OpenStudy (melissa_something):

@bibby

OpenStudy (melissa_something):

OMG NVM I got it that was really easy. It's always these that get me.

OpenStudy (bibby):

what was it?

OpenStudy (melissa_something):

the same thing but squared ^2 cancels the radical thats the only thing that changed

OpenStudy (bibby):

correct (x-7)^2+(y--4)^2=19

OpenStudy (anonymous):

... Medal*...